Question 3 Two variables are defined, a regression equation is given, and one data point is given. Study = number of hours spent studying for an exam Grade = grade on the exam Grade 41.13.5(Study) The data point is a student who studied 10 hours and received an 81 on the exam. (a) Find the predicted value for the data point and compute the residual. Enter the exact answers. Predicted value = Residual =
Answer #1

Predicted value = 41.1+3.5*10 = 76.1

Residual = Actual - Predicted = 81-76.1 = 4.9